Newtons laws of cooling proposes that the rate of change of temperature is proportional to the temperature difference to the ambient (room) temperature. And can be modelled using the equation: dT/dt = -k (T-Ta)

It can also be written as dT/T-Ta = -k dt

Where:
T = Temperature of material
Ta = Ambient (room) temperature
k = A cooling constant

a) integrate both sides of the equation and show that the temperature difference is given by:

(T-Ta) = CoE^-kt

(Co is a constant for this problem)

B) calculate Co if the initial temperature is 70 degrees C and Ta = 20 degrees C?
1
Expert's answer
2019-11-27T10:30:23-0500

A) 

"\int \cfrac{dT}{T-T_a} = - \int k\,dt"


"ln\,|T - T_a| = -kt + ln\,C_0"


"T - T_a = e^{-kt + ln\,C_0}"


"T - T_a = e^{ln\,C_0}\,e^{-kt}"


"T - T_a = C_0\,e^{-kt}"


B)

"70 - 20 = C_0\,e^{-k*0} \\
\underline {C_0 = 50}"